Care and Feeding of the 48 inch Telescope 03/23/94

The 48” telescope is an f/8 Ritchey-Chretien reflector, using a honeycomb borosilicate f/1.9 primary mirror and giving a coma free field of 30 arcminutes. There are two secondary mirrors, each giving the same cassegrain focal ratio, but one being undersized for infrared observations. The primary mirror is supported in its cell by an air piston system, using two o-rings at the base of the mirror to retain the air pressure. The axial definition is by three pads, which have load sensors to set the air pressure. The mirror is defined radially via a central hub, which is in turn pinned to the mirror cell hub via 4 bolts. There are a number of safety features to ensure that the mirror is not overpressurized. The two most notable are the pressure controller in the telescope rack in the control room which will click quietly every now and then, and a rapid vent underneath the mirror cell, which will make loud noises whenever the telescope is slewed over large zenith distance differences.

The secondaries are centrally supported and defined. The secondary mechanism (“pod”) potentially allows full remote control of the position of the mirror, however in reality it has proved to be such a source of problems, that only focus movement is routinely used.

The fork mount uses direct friction drives, allowing very fast slewing but having the potentially dangerous feature that the telescope can be moved by hand when the motors are not engaged. The motors are controlled via a MacIntosh computer running software supplied by the builder, DFM Engineering. The dome position is encoded, and when properly set will follow the motion of the telescope.

An autoguider program running on a PC allows movement of the TV pickoff mirror, in the Top Box, focus of the TV, as well as autoguiding. The TV field is presently about 2 arcminutes. The Top Box also has an 8 position filter wheel, which is controlled by the Sun computer. The maximum filter size is 4”, inserts allow 2” square and round filters also to be used. Filters are changeable only by qualified people.

(I) Startup If the coordinates are good or if the power has been left on, skip to (d). The DFM rack has numbers placed next to the switches, indicating the order in which the switches are to be turned on.

(a) Switch on the power (1)

(b) Push the Mac boot button (2). When the APX cursor appears, type 3:tcs A menu will appear. Some important things to know are: RETURN

backs up to the next higher menu, ESC aborts a menu command. The rest is obvious. The program does not automatically read the Mac clock, so to do that, go into the Initialization Menu, enter into the SET DATE and TIME menu, and hit returns until the clock is read. Now back up to the Main Menu. We are still having trouble with both the Mac clock and the DFM clock, so it seems to be a good idea to check the clock once a day, and reset the DFM clock with the Mac clock.

(c) Turn on the telescope power (3).

(Start here for normal operation)

(d) Turn on the Aux Rack power (4), and the pressure meter power (5).

(e) Check the Ammeter to see that the gauge reads about 0.64 mAmp. If the reading is larger than that, call Bas or see his notes on the primary mirror support.

(f) Read manometer pressure (on top of the DFM rack), it should be zero.

(g) Switch on the primary air pump (6). The air pressure should rise on the manometer to between 0.18 and 0.19.

(h) Switch on the dome rotation power (7)

(i) Switch on the drive power (8)

(j) and the drive motors (9). (the button is always red, when the button is out the drives are on).

If the telescope coordinates are ok (things weren’t turned off from the last night). then skip to (l).

(k) Use the levels on the telescope to align the telescope to the zenith. Now go to the Mac terminal, enter the Initialization Menu, and set the telescope position. Since the software requires entering an RA rather than an HA, enter the present ST as RA, the Dec should be 31 41 17, current epoch. Wait till the ST rolls around before hitting the final Return.

(l) Now, if the sky is ready, open the dome slit. Be sure to unplug the power cord

(m) Open the mirror doors (10). It’s important to open the dome slit first so that dirt and such do not fall on the mirror. Do not open the mirror doors unless the telescope is at the zenith, otherwise the doors may get stuck.

(n) Turn on the tracking. Move to a bright star, by going to the Movement Menu and using the Set Slew command. That command loads the coordinates only, you have to type 7 now (or push the GO button) to actually move the telescope. If necessary, check the finder telescopes to locate the star. Turn on the TV power and High Voltage out underneath the telescope. In the control room, turn on the two TV monitors and the cursor box. Move the Top Box mirror to the axis position, turn up the TV gain with the pot, and center the star. On the Mac, go back into the Initialization Menu and set the telescope position, you may be able to enter zeros if the star’s coords are still listed as Next Object.

(o) Go out into the dome and move the dome with the hand paddle until the telescope points out. On the Mac, go back into the Initialization Menu and set the dome azimuth equal to telescope azimuth. Turn on the auto dome switch (11), and the dome track switch (12).

You should be set to observe now. Because there appears to be some slippage in the encoders it is advisable to occasionally reset the coordinates with a bright star, particularly if you have gone to large zenith distances.

(II) Miscellaneous things

The track rate should be set to 15.09 for best performance as of Feb 92. To do this go to the Rates Menu, enter the track rate menu, set the RA rate to 15.09.

The handpaddle rates can be set in the same Menu, in particular the guide rate should be set to around 0.3 for the autoguider to work. Also, make sure the External Computer switch is on when you are autoguiding (it can be left on at all times).

The telescope can now be slewed from the Sun, however when the Mac communication is on, commands entered into the Mac directly are ignored every other time. So, if you want to enter commands into the Mac, turn off the Mac Communication in the TCS window on the Sun, or enter the commands twice, usually the second time only involves hitting a bunch of Returns.

The Display Epoch can be arbitrarily changed in the Miscellaneous Menu. You can enter coords into a table in that menu as well. This is the epoch that will be stored in the data header.

The Year displayed on the Mac is rounded off, so after June, the displayed year will read current year + 1.

There is an evaporative cooler in the dome which should be used to keep the dome as cool as possible. The on switch is in the fuse box between the inside and outside doors.

(III) Flats

Open the mirror doors. Move the dome such that the white spot is due north. Slew the telescope by hand to about 84 degrees North, 0 HA (tracking off). Check the positioning by eye. Turn on the lamps with the switch on the DFM rack, make sure all lights are on.

There are several macros in the Realtime system that allow you to take flats in all of your filters with a single command, say while you are eating dinner. Be sure to turn off the lamps when you are done.

(IV) Shutdown

Basically, you reverse the order of Startup, but leave on switches 1, 2 and 3.

(a) Move the telescope to the zenith by using the 5 command in the Movement Menu. It will remind you to turn off the tracking first. You also have to type 7 to execute the move.

(b) Switch the dome track to dome home, when the dome is home, turn off auto dome (11).

(c) Close the mirror doors (10).

(d) Now switch off in this order, switches 9, 8, 7, 6, 5, and 4. Also turn off all monitors.

(e) Close the dome, the lower shutter should be closed before the upper one makes contact.

(f) Switch off the TV in the dome.

(V). PROBLEMS

The DFM software has a protection against slewing too far over, thus the slew command will fail with the message ”OUT OF RANGE”. It is possible however to go right up to the limit with the slew command, and perhaps encounter problems, like ”APPROACHING LIMIT”. Also, it is possible to go to the limit by slewing with the hand paddle.

If you have reached the physical limit of 20 degrees above the horizon, the motors will shut off automatically. To get restarted, first and most important, find out what you did wrong and correct that. It occassionally happens that you have reset the coordinates completely wrong, and then tried to slew to what would have been a reasonable place, but which was rendered impossible to achieve by the zero point mistake. Once you know what you did wrong, go out to the telescope and gently push on the top ring in an upwards direction until the motors re-engage. You can now slew the telescope by hand till it is in a reasonable position.

The dome occasionally gets lost. To retrieve the pointing of the dome, turn off the autodome switch on the DFM rack. Go out into the dome and rotate it with the hand paddle until the telescope is pointing out the middle of the slit. On the Mac computer, go to the initialization menu, and type 3 for the dome initialization command. Enter 0 when prompted, thus making the dome azimuth equal to the telescope azimuth. Check the display to make sure the command was accepted, since the Sun-Mac communication often causes commands to be lost. Now turn the autodome back on.